OpenStudy (niahloves):

Select the expression that is equivalent to 3(n + 6). A. 2n + 2 + n + 4 B. 3n + 18 C. 3n + 6 D. 2(n + 6) + n

OpenStudy (niahloves):

@YK.PAPI

OpenStudy (anonymous):

wat do u think

OpenStudy (niahloves):

idk

rishavraj (rishavraj):

hint: a(x + y) = ax + ay

OpenStudy (anonymous):

thanks i was jus gonna type dat

OpenStudy (niahloves):

C????

OpenStudy (anonymous):

nice gud job

rishavraj (rishavraj):

nah nah its not C

rishavraj (rishavraj):

3(n + 6) = 3n + 3*6 = 3n + 18

rishavraj (rishavraj):

@niahloves